OpenStudy (anonymous):

why does 100! have 24 zeroes

ganeshie8 (ganeshie8):

prime factorize : 100! = 2^m * 3^n * 5^r.... to produce a 0 at the end, you will need a 2 and a 5 since there will be plenty of 2's compared to 5's, we need to check oly how many 5's area available to multiply wid 2

ganeshie8 (ganeshie8):

number of 5's = 100/5 + 100/5^2 = 20 + 4 = 24

ganeshie8 (ganeshie8):

prime factorization will have : 5^24 and 2 power will be larger than that. so there will be exactly 24 0's in the tail of in 100!

OpenStudy (anonymous):

My problem is to approach it with a list view of the numbers. Like this. The number of zeroes found in the ten's - that's 11 because there are two zeroes in 100 The number of 5's found in the 5's - that's ten The number of 5's found in 5^2 numbers - that's 3 Which comes to the same result. Thank you for a more rigorous approach

ganeshie8 (ganeshie8):

looks good :) both methods are one and same actually !